If the soda can was at a pressure of 160.3 kPa at 393K, what was the pressure at 258K when it imploded if the volume was the same before and after?


1
Expert's answer
2021-01-15T06:39:16-0500

"\frac{P_1}{T_1}" = "\frac{P_2}{T_2}" , P1 = 160.3kPa , T1 = 393K , P2 = ? , T2 = 258K

"\frac{160.3}{393}" = "\frac{P_2}{258}"

P2 = "\frac{258×160.3}{393}"

P2 = 105.23kPa